# MSPCaps: A Multi-Scale Patchify Capsule Network with Cross-Agreement Routing for Visual Recognition

### v1

@@ Abstract
Capsule Network (CapsNet) has demonstrated significant potential in visual recognition by capturing spatial relationships and part-whole hierarchies for learning equivariant feature representations. However, existing CapsNet and variants often rely on a single high-level feature map, overlooking the rich complementary information provided by multi-scale features. Furthermore, conventional feature fusion strategies, such as addition and concatenation, struggle to reconcile multi-scale feature discrepancies, leading to suboptimal classification performance. To address these limitations, we propose the Multi-Scale Patchify Capsule Network (MSPCaps), a novel architecture that integrates multi-scale feature learning and efficient capsule routing. Specifically, MSPCaps consists of three key components: a Multi-Scale ResNet Backbone (MSRB), a Patchify Capsule Layer (PatchifyCaps), and a Cross-Agreement Routing (CAR) block. First, the MSRB extracts diverse multi-scale feature representations from input images, preserving both fine-grained details and global contextual information. Second, the PatchifyCaps partitions these multi-scale features into primary capsules using a uniform patch size, equipping the model with the ability to learn from diverse receptive fields. Finally, the CAR block adaptively routes the multi-scale capsules by identifying cross-scale prediction pairs with maximum agreement. Unlike the simple concatenation of multiple self-routing blocks, CAR ensures that only the most coherent capsules (best part-to-whole pairs) contribute to the final voting. Our proposed MSPCaps achieves remarkable scalability and superior robustness, consistently surpassing multiple baseline methods in terms of classification accuracy, with configurations ranging from a highly efficient Tiny model (344.3K parameters) to a powerful Large model (10.9M parameters), highlighting its potential in advancing feature representation learning.
11footnotetext: Corresponding Author.
@@ Introduction
Convolutional Neural Networks (CNNs) have proven exceptional capabilities in extracting hierarchical image representations. However, understanding visual scenes requires not only detecting parts or objects, but also understanding their spatial relationships and part-whole structures. The inability of CNNs to inherently model these relationships leads to significant limitations, including low robustness and poor spatial object correlation. To address these issues, Capsule Network (CapsNet), first introduced by [23], encodes entities as vectors representing intrinsic properties, such as pose and orientation. In simple terms, CNNs determine if a feature is present, while CapsNet explains how it exists by capturing the relationships between parts.
@@
Recent studies have advanced capsules with deep backbones [22, 14, 5] or attention routing [19, 1, 21]. However, despite their compelling designs, CapsNet and its variants still face several limitations: (1) Existing studies [19] instantiate primary capsules directly from entire global feature maps. This runs counter to the intuition that the lowest-level capsule should focus on simple visual primitives that are spatially localized; (2) Existing capsule models mainly rely on single-resolution features, neglecting the rich complementary information available across different scales; (3) Conventional feature fusion strategies, such as simple addition and concatenation, struggle to effectively reconcile the discrepancies between multi-scale features, thereby falling short in leveraging multi-scale features.
To address these limitations, we propose the Multi-Scale Patchify Capsule Network (MSPCaps), a novel architecture that integrates multi-scale feature learning and efficient capsule routing. Fig. 1 presents a comparative overview of the conventional CapsNet and our MSPCaps model. Specifically, MSPCaps consists of three core components: a Multi-Scale ResNet Backbone (MSRB), a Patchify Capsule Layer (PatchifyCaps), and a Cross-Agreement Routing (CAR) block. Firstly, MSRB utilizes a shared ResNet backbone to extract a hierarchy of multi-scale features from input images. Typically, images with different resolutions contain complementary information: high-resolution features tend to preserve fine-grained details (local primitives), while low-resolution representations provide global semantic information (object-level entities). Based on this, PatchifyCaps takes the feature maps from different scales and partitions each feature map into a grid of fixed-sized, non-overlapping patches, generating spatially localized primary capsules. This mechanism not only enables the model to focus on local regions instead of global features but also allows it to learn from diverse receptive fields, which capture intricate patterns at high resolutions and broader abstractions at lower resolutions. Finally, to effectively fuse multi-scale features, the CAR block aggregates multi-scale capsules by identifying cross-scale prediction pairs with maximum agreement, thereby ensuring that only the most coherent capsules (best part-to-whole pairs) contribute to the final decision. Moreover, by progressively fusing high and low resolution capsules, CAR provides rich guidance for generating complex higher-level capsules and acts as an implicit regularizer, preventing the model from being misled by excessive fine-grained details. The effectiveness of the proposed MSPCaps is evaluated on four baseline datasets: MNIST, FashionMNIST, CIFAR-10, and SVHN. The results show that MSPCaps achieve state-of-the-art performance and superior robustness compared to multiple baselines.
@@
Routing Techniques in Capsule Networks. Research on CapNets has explored various routing strategies to improve efficiency and representation capabilities. Early methods such as Dynamic Routing (DR) [23] and EM Routing [12] employ iterative clustering mechanisms to refine coupling coefficients, using vector or matrix representations, respectively. More recently, attention-based [1, 14] and non-iterative [2, 6] strategies have gained prominence. Specifically, Efficient-Caps [19] introduces a self-attention routing scheme to compute coupling coefficients based on similarity among capsule votes, enabling fully parallel and non-iterative routing. OrthCaps [5] further enhances efficiency by integrating pruning and orthogonal sparse attention routing, effectively reducing redundancy while maintaining accuracy. Moreover, SR-CapsNet [9] draws inspiration from the Mixture-of-Experts (MOE) architecture, using two sub-networks per capsule to generate pose vectors and activations, achieving expert-style routing. However, those methods predominantly rely on single-resolution features, and little attention has been paid to multi-scale routing strategies.
Multi-Scale Feature and Fusion. Multi-scale feature modeling has been widely explored to improve visual classification by capturing both local details and global context. CNN-based methods [18, 26, 4, 13] incorporate multi-scale information and fuse them following a common pattern during fusion: combining the features via concatenation or addition, followed by projection through convolutions. Similarly, CapsNet variants such as MS-CapsNet [27] and RS-CapsNet [29] fuse capsules across scales by concatenating and jointly routing multi-scale capsules, following a comparable fusion paradigm. Nevertheless, the multi-scale fusion in CapNets remains shallow, as most existing approaches concatenate or add features across scales without deeper interaction, limiting the effectiveness of feature integration.
Figure 2: (a) The overall architecture of MSPCaps, which consists of an MSRB, a PatchifyCaps, and two CAR blocks. The backbone is designed with three residual blocks to generate features at varying scales, which are then processed by the PatchifyCaps and CAR blocks. (b) The structure of each residual block, which starts with a downsampling convolution (stride=2=2) followed by BatchNorm and ReLU. The output is then succeeded by $n$ identical convolutional modules, with each one a residual connection.
To address the aforementioned limitations, we propose MSPCaps that leverage multi-scale features to obtain diverse receptive fields and progressively fuse information across scales through a novel routing mechanism, leading to more effective feature integration and improved performance.
@@
We propose MSPCaps, a novel architecture built upon a Multi-Scale ResNet Backbone (MSRB) that captures both fine-grained details and global semantic information. This architecture integrates two key components: (1) a Patchify Capsule Layer (PatchifyCaps), which learns complementary information across different receptive fields and transforms these features into primary capsules; and (2) a Cross-Agreement Routing (CAR) block for adaptive cross-scale capsule voting. To demonstrate the scalability of the proposed architecture, we present two model versions: a lightweight model–MSPCaps-T and a powerful model–MSPCaps-L. Both models share the same core design and differ only in key hyperparameters and the utilization of the weight-sharing mechanism within CAR.
@@
The MSRB is designed to generate hierarchical features. Inspired by the progressive downsampling strategy in ResNet [11], MSRB first generates the finest feature map $u_{1}$ and progressively reduces spatial resolution using convolutions with a kernel size of 3 and a stride of 2. Deeper feature learning is achieved using subsequent convolutions with a kernel size of 3 and a stride of 1.
The primary motivation for this design is to control the learning process at different levels of abstraction and align it with the subsequent PatchifyCaps. For high-resolution (fine-grained) feature maps, we employ a shallower network with fewer parameters, encouraging the model to efficiently capture essential local details while reducing the risk of overfitting caused by overly complex fine-grained features. Conversely, for low-resolution (coarse-grained) feature maps, which encapsulate global context, we utilize a deeper network with richer parameterization to enhance abstract representations learning. Furthermore, this structure introduces a potent regularization effect that stems from the use of a hierarchical weight-sharing mechanism. Specifically, the initial residual blocks are shared across all scales, while deeper blocks are specialized for coarse resolutions. This dual supervision updates the shared kernels with gradients from both the fine and coarse outputs, which implicitly regularizes the network to extract more general and robust representations.
@@
A key challenge in multi-scale fusion lies in preserving both critical fine-grained and global information. To address this, we propose PatchifyCaps, a module that interfaces with MSRB to generate multi-scale capsules. Inspired by patch-based architectures such as Vision Transformers [3], PatchifyCaps partitions multi-scale feature maps into a grid of patches. By applying a uniform patch size to feature maps of different scales, the capsules derived from high-resolution maps capture local patterns via small receptive fields, while those from coarser maps gain a wider view to model global structures. Beyond creating primary capsules, PatchifyCaps intrinsically links receptive field size to the feature hierarchy, enabling adaptive perception across scales.
To implement this, PatchifyCaps employs a “one-patch, one-capsule" mapping. It first partitions each feature map into a grid of non-overlapping patches using an average pooling layer. This partitioning strategy enforces a strict local receptive field, ensuring each resulting capsule is sensitive to spatial patterns within its designated region. Subsequently, a $1\times 1$ convolution adapts the feature channels to the desired capsule dimension, $d_{i}$. The resulting tensor is then reshaped into a sequence of capsules, $u_{i}\in\mathbb{R}^{n_{i}\times d_{i}}$. To compensate for the spatial context lost during the flattening step, we augment each capsule vector with a learnable positional embedding. This step ensures that each capsule retains its region-sensitive properties, encoding both the presence and location of features on the feature map grid. The detailed calculation process is summarized in Alg. 1.
@@
The “one-patch, one-capsule" approach is computationally efficient. By replacing parallel convolutions with a simple pooling layer and a single $1\times 1$ convolution, it significantly reduces the number of parameters required to form the primary capsules. Moreover, we can effectively compress the number of primary capsules through an appropriate patch size. For instance, when $p=4$, the number of primary capsules is reduced to just 84 compared to 1152 in [23], with a decrease of over 90%. This drastic reduction proportionally decreases the parameters and computations needed for weight projection during voting, yielding a substantial downstream effect.
@@
The efficacy of conventional capsule routing is often limited by its single-source design, which fails to leverage the rich, multi-scale information essential for complex feature voting. Although prior works like RS-CapsNet [29] incorporate capsules from various scales, their fusion remains simplistic, merely aggregating multiple intermediate capsules for final routing. To overcome this limitation, we introduce CAR, a novel routing mechanism to achieve sophisticated agreement-based voting across feature scales. As shown in Fig. 3, the CAR begins with two parallel input capsules from a lower layer $l$: finer capsules $u_{1}^{l}\in\mathbb{R}^{n_{\text{in}}^{(1)}\times d_{\text{in}}^{(1)}}$ representing smaller objects, and coarser capsules $u_{2}^{l}\in\mathbb{R}^{n_{\text{in}}^{(2)}\times d_{\text{in}}^{(2)}}$ containing larger ones. Both capsules are first transformed into prediction capsules $\hat{u}^{(1)}$ and $\hat{u}^{(2)}$ for the $n_{\text{out}}$ output capsules in layer $l+1$. This process is achieved by applying transformation matrices, $W^{(1)}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}^{(1)}\times d_{\text{in}}^{(1)}\times d_{\text{out}}}$ and $W^{(2)}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}^{(2)}\times d_{\text{in}}^{(2)}\times d_{\text{out}}}$, which can be given by
@@
2:if share weight then
@@
The foundation of our routing mechanism is based on a key hypothesis: The uniform patch size for cross-scale capsules ensures inherent spatial correspondence. Specifically, a local group of $s=n_{\text{in}}^{(1)}/n_{\text{in}}^{(2)}$ capsules from the finer input $u_{1}^{l}$ corresponds spatially to a single capsule from the coarser input $u_{2}^{l}$, resembling the parts-to-whole relationships in traditional dynamic routing. Therefore, capsules within the same local region should exhibit high similarity and thus make a consistent contribution for voting. Building on this premise, we introduce a lightweight design (weight-sharing mechanism) to reduce parameters for MSPCaps-T by assuming that the mutual information between two scales of capsules is sufficient for classification. To implement this, we simply apply transformation matrix $W^{(2)}$ to the finer capsules $u_{1}^{l}$ to generate prediction capsules $\hat{u}^{(1)}$, which is given by
@@
where $\{\hat{u}_{j|m,k}^{(1)}\}_{m=1}^{s}$ is the group of prediction vectors from the finer capsules that spatially correspond to the $k$-th coarser capsule. By adaptively selecting the maximum similarity, the most representative feature alignment (best part-to-whole pair) within the local patch guides the routing process. The raw agreement scores are then normalized using the softmax function to compute the final coupling coefficients, $c_{jk}$. The resulting output capsule $v_{j}$ is then calculated as a weighted sum over the coarser-scale prediction vectors, $v_{j}=\sum_{k}c_{jk}\hat{u}_{j|k}^{(2)}$, which is then passed through a squash activation function [23] to generate the intermediate capsules or class capsules.
@@
In our experiments, we select four baseline datasets: MNIST [17], FashionMNIST [28], SVHN [20], and CIFAR-10 [15], which are widely used in the literature of CapsNet. During data preprocessing, the data samples in MNIST and Fashion-MNIST are resized from $28\times 28$ to $32\times 32$, enabling performance comparison under the same model configuration. To enhance generalization, we apply data augmentation to the training splits only. Specifically, the images in Fashion-MNIST, SVHN, and CIFAR-10 undergo a $32\times 32$ random crop with 4-pixel padding on each side, followed by a random horizontal flip except for SVHN. For the MNIST and SVHN datasets, we apply random rotations within $\pm 15$ degrees.
@@
All experiments are conducted using PyTorch version 2.5.1 with Python 3.12 and CUDA 12.4 on a single NVIDIA RTX 4090D GPU. In particular, we utilize the margin loss as defined in [23]. The reconstruction loss is omitted as it provides only a negligible improvement in model performance. We train our model for 300 epochs using a batch size of $128$. Additionally, we employ the AdamW optimizer with a learning rate of $5\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. We also use a learning rate schedule with a 5-epoch linear warmup from 10% of the base value, followed by a cosine annealing decay for the remaining $295$ epochs to a minimum of $1\times 10^{-6}$. Other hyperparameters of our model variants are provided in Tab. 1.
Hyperparameters MSPCaps-T MSPCaps-L
@@
To demonstrate the advantage of the proposed MSPCaps, we compare it with 9 state-of-the-art models, including: (1) CapsNet [23]; (2) AA-CapsNet [21]; (3) AR-CapsNet [1]; (4) DA-CapsNet [14]; (5) DeepCaps [22]; (6) RS-CapsNet [29]; (7) CapsNet with inverted dot-product (IDP-CapsNet) [25]; (8) Efficient-Caps [19]; and (9) OrthCaps [5].
@@
Tab. 2 and Tab. 3 illustrate the performance of MSPCaps-T and MSPCaps-L against the baseline models. The results for each model are based on the best performance achieved during training with random seed initialization. The asterisk (*) indicates that the result is reported directly from the original literature, as official code is unavailable.
@@
Table 2: Performance comparison of MSPCaps-T with other baseline models on 3 image classification datasets. Best results are in bold.
@@
Table 3: Performance comparison of MSPCaps-L with other baseline models on 3 image classification datasets. Best results are in bold. indicates that the result is obtained using an ensemble of 7 models.

### v2

@@ Abstract
Capsule Network (CapsNet) has demonstrated significant potential in visual recognition by capturing spatial relationships and part-whole hierarchies for learning equivariant feature representations. However, existing CapsNet and variants often rely on a single high-level feature map, overlooking the rich complementary information provided by multi-scale features. Furthermore, conventional feature fusion strategies, such as addition and concatenation, struggle to reconcile multi-scale feature discrepancies, leading to suboptimal classification performance. To address these limitations, we propose the Multi-Scale Patchify Capsule Network (MSPCaps), a novel architecture that integrates multi-scale feature learning and efficient capsule routing. Specifically, MSPCaps consists of three key components: a Multi-Scale ResNet Backbone (MSRB), a Patchify Capsule Layer (PatchifyCaps), and a Cross-Agreement Routing (CAR) block. First, the MSRB extracts diverse multi-scale feature representations from input images, preserving both fine-grained details and global contextual information. Second, the PatchifyCaps partitions these multi-scale features into primary capsules using a uniform patch size, equipping the model with the ability to learn from diverse receptive fields. Finally, the CAR block adaptively routes the multi-scale capsules by identifying cross-scale prediction pairs with maximum agreement. Unlike the simple concatenation of multiple self-routing blocks, CAR ensures that only the most coherent capsules (best part-to-whole pairs) contribute to the final voting. Our proposed MSPCaps achieves remarkable scalability and superior robustness, consistently surpassing multiple baseline methods in terms of classification accuracy, with configurations ranging from a highly efficient Tiny model (344.3K parameters) to a powerful Large model (10.9M parameters), highlighting its potential in advancing feature representation learning. The code is available at https://github.com/abdn-hyd/MSPCaps.
@@ Introduction
Convolutional Neural Networks (CNNs) have proven exceptional capabilities in extracting hierarchical image representations. However, understanding visual scenes requires not only detecting parts or objects, but also understanding their spatial relationships and part-whole structures. The inability of CNNs to inherently model these relationships leads to significant limitations, including low robustness and poor spatial object correlation. To address these issues, Capsule Network (CapsNet), first introduced by (Sabour et al. 2017), encodes entities as vectors representing intrinsic properties, such as pose and orientation. In simple terms, CNNs determine if a feature is present, while CapsNet explains how it exists by capturing the relationships between parts.
@@
Recent studies have advanced capsules with deep backbones (Rajasegaran et al. 2019; Huang and Zhou 2020; Geng et al. 2024) or attention routing (Mazzia et al. 2021; Choi et al. 2019; Pucci et al. 2021). However, despite their compelling designs, CapsNet and its variants still face several limitations: (1) Existing studies (Mazzia et al. 2021) instantiate primary capsules directly from entire global feature maps. This runs counter to the intuition that the lowest-level capsule should focus on simple visual primitives that are spatially localized; (2) Existing capsule models mainly rely on single-resolution features, neglecting the rich complementary information available across different scales; (3) Conventional feature fusion strategies, such as simple addition and concatenation, struggle to effectively reconcile the discrepancies between multi-scale features.
To address these limitations, we propose the Multi-Scale Patchify Capsule Network (MSPCaps), a novel architecture that integrates multi-scale feature learning and efficient capsule routing. Fig. 1 presents a comparative overview of the conventional CapsNet and our MSPCaps model. Specifically, MSPCaps consists of three core components: a Multi-Scale ResNet Backbone (MSRB), a Patchify Capsule Layer (PatchifyCaps), and a Cross-Agreement Routing (CAR) block. Firstly, MSRB utilizes a shared ResNet backbone to extract a hierarchy of multi-scale features from input images. Typically, images with different resolutions contain complementary information: high-resolution features tend to preserve fine-grained details (local primitives), while low-resolution representations provide global semantic information (object-level entities). Based on this, PatchifyCaps takes the feature maps from different scales and partitions each feature map into a grid of fixed-sized, non-overlapping patches, generating spatially localized primary capsules. This mechanism not only enables the model to focus on local regions but also allows it to learn from diverse receptive fields, which capture intricate patterns at high resolutions and broader abstractions at lower resolutions. Finally, to effectively fuse multi-scale features, the CAR block aggregates multi-scale capsules by identifying cross-scale prediction pairs with maximum agreement, thereby ensuring that only the most coherent capsules (best part-to-whole pairs) contribute to the final decision. Moreover, by progressively fusing high and low resolution capsules, CAR provides rich guidance for generating complex higher-level capsules and acts as an implicit regularizer, preventing the model from being misled by excessive fine-grained details. The effectiveness of the proposed MSPCaps is evaluated on four baseline datasets: MNIST, FashionMNIST, CIFAR-10, and SVHN. The results show that MSPCaps achieve state-of-the-art performance and superior robustness compared to multiple baselines within capsule networks.
@@
Figure 2: (a) The overall architecture of MSPCaps, which consists of an MSRB, a PatchifyCaps, and two CAR blocks. The backbone is designed with three residual blocks to generate features at varying scales. (b) Each residual block starts with a downsampling convolution (stride=2), followed by $n$ identical convolutional modules, each with a residual connection.
Routing Techniques in Capsule Networks. Research on CapNets has explored various routing strategies to improve efficiency and representation capabilities. Early methods such as Dynamic Routing (DR) (Sabour et al. 2017) and EM Routing (Hinton et al. 2018) employ iterative clustering mechanisms to refine coupling coefficients, using vector or matrix representations, respectively. More recently, attention-based (Choi et al. 2019; Huang and Zhou 2020) and non-iterative (Ding et al. 2020; Geng et al. 2025) strategies have gained prominence. Specifically, Efficient-Caps (Mazzia et al. 2021) introduces a self-attention routing scheme to compute coupling coefficients based on similarity among capsule votes, enabling fully parallel and non-iterative routing. OrthCaps (Geng et al. 2024) further enhances efficiency by integrating pruning and orthogonal sparse attention routing, effectively reducing redundancy while maintaining accuracy. Moreover, SR-CapsNet (Hahn et al. 2019) draws inspiration from the Mixture-of-Experts (MOE) architecture, using two sub-networks per capsule to generate pose vectors and activations, achieving expert-style routing. However, those methods predominantly rely on single-resolution features, and little attention has been paid to multi-scale routing strategies.
Multi-Scale Feature and Fusion. Multi-scale feature modeling has been widely explored to improve visual classification by capturing both local details and global context. CNN-based methods (Lin et al. 2017; Wang et al. 2020; Gao et al. 2019; Huang et al. 2017) incorporate multi-scale information and fuse them following a common pattern during fusion: combining the features via concatenation or addition, followed by projection through convolutions. Similarly, CapsNet variants such as MS-CapsNet (Xiang et al. 2018) and RS-CapsNet (Yang et al. 2020) fuse capsules across scales by concatenating and jointly routing multi-scale capsules, following a comparable fusion paradigm. Nevertheless, the multi-scale fusion in CapNets remains shallow, as most existing approaches concatenate or add features across scales without deeper interaction, limiting the effectiveness of feature integration.
@@
We propose MSPCaps, a novel architecture constructed from three core modules: (1) a Multi-Scale ResNet Backbone (MSRB), designed to capture both fine-grained details and global semantic information; (2) a Patchify Capsule Layer (PatchifyCaps), which learns complementary information across different receptive fields and transforms the extracted features into primary capsules; (3) and a Cross-Agreement Routing (CAR) block for adaptive cross-scale capsule voting. To demonstrate the scalability of the proposed architecture, we present two model versions: a lightweight model–MSPCaps-T and a powerful model–MSPCaps-L. Both models share the same core design and differ only in key hyperparameters and the utilization of the weight-sharing mechanism within CAR.
@@
The MSRB is designed to generate hierarchical features. Inspired by the progressive downsampling strategy in ResNet (He et al. 2016), MSRB first generates the finest feature map $u_{1}$ and progressively reduces spatial resolution using convolutions with a kernel size of 3 and a stride of 2. Deeper feature learning is achieved using subsequent convolutions with a kernel size of 3 and a stride of 1.
The primary motivation for this design is to control the learning process at different levels of abstraction and align it with the subsequent PatchifyCaps. For high-resolution (fine-grained) feature maps, we employ a shallower network with fewer parameters, encouraging the model to capture essential local details while reducing the risk of overfitting caused by overly complex fine-grained features. Conversely, for low-resolution (coarse-grained) feature maps, which encapsulate global context, we utilize a deeper network with richer parameterization to enhance abstract representations learning. Furthermore, this structure introduces a potent regularization effect that stems from the use of a hierarchical weight-sharing mechanism. Specifically, the initial residual blocks are shared across all scales, while deeper blocks are specialized for coarse resolutions. This dual supervision updates the shared kernels with gradients from both the fine and coarse outputs, which implicitly regularizes the network to extract more general and robust representations.
@@
A key challenge in multi-scale feature representation lies in preserving both critical fine-grained and global information. To address this, we propose PatchifyCaps, a module that interfaces with MSRB to generate multi-scale capsules. Inspired by patch-based architectures such as Vision Transformers (Dosovitskiy et al. 2020), PatchifyCaps partitions multi-scale feature maps into a grid of patches. By applying a uniform patch size to feature maps of different scales, the capsules derived from high-resolution maps capture local patterns via small receptive fields, while those from coarser maps gain a wider view to model global structures. Beyond creating primary capsules, PatchifyCaps intrinsically links receptive field size to the feature hierarchy, enabling adaptive perception across scales.
To implement this, PatchifyCaps employs a “one-patch, one-capsule” mapping. It first partitions each feature map into a grid of non-overlapping patches using an average pooling layer. This partitioning strategy enforces a strict local receptive field, ensuring each resulting capsule is sensitive to spatial patterns within its designated region. Subsequently, a $1\times 1$ convolution adapts the feature channels to the desired capsule dimension, $d_{i}$. The resulting tensor is then reshaped into a sequence of capsules, $u_{i}\in\mathbb{R}^{n_{i}\times d_{i}}$. To compensate for the spatial context lost during the flattening step, we augment each capsule vector with a learnable positional embedding. This step ensures that each capsule retains its region-sensitive properties, encoding both the presence and location of features on the feature map grid. The detailed calculation process is summarized in Alg. 1.
@@
The “one-patch, one-capsule” approach is computationally efficient. By replacing parallel convolutions with a simple pooling layer and a single $1\times 1$ convolution, it significantly reduces the number of parameters required to form the primary capsules. Moreover, we can effectively compress the number of primary capsules through an appropriate patch size. For instance, when $p=4$, the number of primary capsules is reduced to just 84 compared to 1152 in (Sabour et al. 2017), with a decrease of over 90%. This drastic reduction proportionally decreases the parameters and computations needed for weight projection during voting, yielding a substantial downstream effect.
@@
The efficacy of conventional capsule routing is often limited by its single-source design, which fails to leverage the rich, multi-scale information essential for complex feature voting. Although prior works like RS-CapsNet (Yang et al. 2020) incorporate capsules from various scales, their fusion remains simplistic, merely aggregating multiple intermediate capsules for final routing. To overcome this limitation, we introduce CAR, a novel routing mechanism to achieve sophisticated agreement-based voting across feature scales. As shown in Fig. 3, the CAR begins with two parallel input capsules from a lower layer $l$: finer capsules $u_{1}^{l}\in\mathbb{R}^{n_{\text{in}}^{(1)}\times d_{\text{in}}^{(1)}}$ representing smaller objects, and coarser capsules $u_{2}^{l}\in\mathbb{R}^{n_{\text{in}}^{(2)}\times d_{\text{in}}^{(2)}}$ containing larger ones. Both capsules are first transformed into prediction capsules $\hat{u}^{(1)}$ and $\hat{u}^{(2)}$ for the $n_{\text{out}}$ output capsules in layer $l+1$. This process is achieved by applying transformation matrices, $W^{(1)}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}^{(1)}\times d_{\text{in}}^{(1)}\times d_{\text{out}}}$ and $W^{(2)}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{in}}^{(2)}\times d_{\text{in}}^{(2)}\times d_{\text{out}}}$, which can be given by
@@
2:if shared weight then
@@
The foundation of our routing mechanism is based on a key hypothesis: The uniform patch size for cross-scale capsules ensures inherent spatial correspondence. Specifically, a local group of $s=n_{\text{in}}^{(1)}/n_{\text{in}}^{(2)}$ capsules from the finer input $u_{1}^{l}$ corresponds spatially to a single capsule from the coarser input $u_{2}^{l}$, resembling the parts-to-whole relationships in traditional dynamic routing. Therefore, capsules within the same local region should exhibit high similarity and thus make a consistent contribution for voting. Building on this premise, we also introduce a lightweight design (weight-sharing mechanism) to reduce parameters for MSPCaps-T by assuming that the mutual information between two scales of capsules is sufficient for classification. To implement this, we simply apply transformation matrix $W^{(2)}$ to the finer capsules $u_{1}^{l}$ to generate prediction capsules $\hat{u}^{(1)}$, which is given by
@@
where $\{\hat{u}_{j|m,k}^{(1)}\}_{m=1}^{s}$ is the group of prediction vectors from the finer capsules that spatially correspond to the $k$-th coarser capsule. By adaptively selecting the maximum similarity, the most representative feature alignment (best part-to-whole pair) within the local patch guides the routing process. The raw agreement scores are then normalized using the softmax function to compute the final coupling coefficients, $c_{jk}$. The resulting output capsule $v_{j}$ is then calculated as a weighted sum over the coarser-scale prediction vectors, $v_{j}=\sum_{k}c_{jk}\hat{u}_{j|k}^{(2)}$, which is then passed through a squash activation function (Sabour et al. 2017) to generate the intermediate capsules or class capsules.
@@
In our experiments, we select four baseline datasets: MNIST (LeCun et al. 1998), FashionMNIST (Xiao et al. 2017), SVHN (Netzer et al. 2011), and CIFAR-10 (Krizhevsky et al. 2009), which are widely used in the literature of capsule networks. During data preprocessing, the data samples in MNIST and Fashion-MNIST are resized from $28\times 28$ to $32\times 32$ for a consistent comparison using the same model configuration. To enhance generalization, we apply data augmentation to the training splits only. Specifically, the images in Fashion-MNIST, SVHN, and CIFAR-10 undergo a $32\times 32$ random crop with 4-pixel padding on each side, followed by a random horizontal flip except for SVHN. For the MNIST and SVHN datasets, we apply random rotations within $\pm 15$ degrees.
@@
All experiments are conducted using PyTorch version 2.5.1 with Python 3.12 and CUDA 12.4 on a single NVIDIA RTX 4090D GPU. In particular, we utilize the margin loss as defined in (Sabour et al. 2017). The reconstruction loss is omitted as it provides only a negligible improvement in model performance. We train our model for 300 epochs using a batch size of $128$. Additionally, we employ the AdamW optimizer with a learning rate of $5\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. We also use a learning rate schedule with a 5-epoch linear warmup from 10% of the base value, followed by a cosine annealing decay for the remaining $295$ epochs to a minimum of $1\times 10^{-6}$. Other hyperparameters of our model variants are provided in Tab. 1.
Hyperparameter MSPCaps-T MSPCaps-L
@@
To demonstrate the advantage of the proposed MSPCaps, we compare it with 9 state-of-the-art models, including: (1) CapsNet (Sabour et al. 2017); (2) AA-CapsNet (Pucci et al. 2021); (3) AR-CapsNet (Choi et al. 2019); (4) DA-CapsNet (Huang and Zhou 2020); (5) DeepCaps (Rajasegaran et al. 2019); (6) RS-CapsNet (Yang et al. 2020); (7) CapsNet with inverted dot-product (IDP-CapsNet) (Tsai et al. 2020); (8) Efficient-Caps (Mazzia et al. 2021); and (9) OrthCaps (Geng et al. 2024).
@@
Tab. 2 and Tab. 3 illustrate the performance of MSPCaps-T and MSPCaps-L against the baseline models. The results for each model are based on the best performance achieved during training with random seed initialization. The asterisk (*) indicates that the result is reported directly from the original paper, as official code is unavailable.
@@
Table 2: Performance comparison of MSPCaps-T with baseline models on MNIST, SVHN and CIFAR-10.
@@
Table 3: Performance comparison of MSPCaps-L with baseline models on MNIST, FashionMNIST and CIFAR-10. indicates the result is obtained using an ensemble of 7 models.